What is the slope of y=2x3 +3x2−3x−2 at x=1?



1
Expert's answer
2022-05-17T08:15:36-0400
"y'=(2x^3+3x^2-3x-2)'"

"=2(3x^2)+3(2x)-3(1)-0"

"=6x^2+6x-3"

"y'(1)=6(1)^2+6(1)-3=9"

The slope of "y=2x^3 +3x^2\u22123x\u22122" at "x=1" is "9."